# GlossLM: Multilingual Pretraining for Low-Resource Interlinear Glossing

### v1

@@ Abstract
A key aspect of language documentation is the creation of annotated text in a format such as interlinear glossed text (IGT), which captures fine-grained morphosyntactic analyses in a morpheme-by-morpheme format. Prior work has explored methods to automatically generate IGT in order to reduce the time cost of language analysis. However, many languages (particularly those requiring preservation) lack sufficient IGT data to train effective models, and crosslingual transfer has been proposed as a method to overcome this limitation.
We compile the largest existing corpus of IGT data from a variety of sources, covering over 450k examples across 1.8k languages, to enable research on crosslingual transfer and IGT generation. Then, we pretrain a large multilingual model on a portion of this corpus, and further finetune it to specific languages. Our model is competitive with state-of-the-art methods for segmented data and large monolingual datasets. Meanwhile, our model outperforms SOTA models on unsegmented text and small corpora by up to 6.6% morpheme accuracy, demonstrating the effectiveness of crosslingual transfer for low-resource languages.111 Our pretrained model and dataset are available at https://huggingface.co/lecslab/glosslm and https://huggingface.co/datasets/lecslab/glosslm-corpus-split, and will be made accessible through a web interface for use in language documentation efforts.
@@
GlossLM: Multilingual Pretraining for Low-Resource Interlinear Glossing
[…]
@@ 1 Introduction
footnotetext: * Equal contribution
With nearly half of the world’s 7,000 languages considered endangered, communities of minoritized language speakers are working to preserve and revitalize their languages (Seifart et al., 2018). These efforts often involve the collection, analysis, and annotation of linguistic data. Annotated text is especially useful in the creation of reference texts (like dictionaries and grammars) and other archival materials, as well as in the development of language technologies including searchable digital text Blokland et al. (2019); Rijhwani et al. (2023) and computer-assisted educational tools Uibo et al. (2017); Chaudhary et al. (2023)
@@
One prevalent form of linguistic annotation in language documentation projects is interlinear glossed text (IGT). IGT is a multi-line data format which includes (1) a transcription of speech in the language, (2) an aligned morpheme-by-morpheme description, and oftentimes (3) a free translation. An example of each of these components is illustrated in Figure 1. IGT is useful for illustrating morphosyntactic features of languages that other linguists or researchers may not be familiar with, and can also be used as examples for grammatical sketches or language learning textbooks, among other applications.
Creating IGT corpora is time-consuming and difficult, requiring documenters to perform linguistic analysis and extensive documentation simultaneously, often for languages with limited prior documentation. Research has found that computational tools can help accelerate the annotation process and overcome this bottleneck Palmer et al. (2009). In particular, models that take as input transcriptions and translations (which don’t require linguistic training to produce) and output glosses can rapidly bootstrap creation of complete IGT datasets.
The majority of prior work on automatic glossing focuses on training monolingual models that can predict IGT for a single language (Moeller and Hulden, 2018; McMillan-Major, 2020; Zhao et al., 2020). This approach has two major drawbacks. First, many language community members or linguists do not have the expertise, time, or resources to create special-purpose models, creating a high barrier of entry for adoption. Second, in the rare cases where expertise is available, there is often insufficient amounts of data to train effective monolingual models, with many IGT corpora having only a few thousand examples or less.
In this work, we aim to overcome the monolingual data bottleneck by creating a multilingual pretrained gloss model that can be further adapted to specific languages and gloss formats. Inspired by recent advances in massively multilingual language modeling, we compile various existing IGT corpora into a dataset containing over 450k IGT examples in 1.8k languages, making it the largest existing collection of digitized IGT. We then train our model, GlossLM, by continually pretraining ByT5 Xue et al. (2022), and evaluate it on seven typologically diverse languages. We find that our model is highly competitive with existing models, outperforming them broadly on the more challenging task of glossing unsegmented text (where words are not split into individual morphemes), as well as in certain cases where segmented text is provided. Furthermore, we observe that our model retains high performance on seen languages even without finetuning, indicating that it can be used as-is without specialization to particular languages.
@@ 2 Interlinear Glossed Text (IGT)
2.1 Description of IGT Format
IGT is a structured data format which presents a transcription of text in a language being studied, along with its gloss—aligned labels that mark each morpheme’s meaning and/or grammatical function—in a different language. Oftentimes, a free translation in a widely-spoken language is included. An IGT example for Arapaho is given in subsection 2.1 Cowell (2020), with glosses and translations in English:
@@ 2.1 Description of IGT Format
This example is segmented, with morphemes separated by dashes. Each morpheme in the Arapaho sentence (e.g. tih) is directly aligned with glosses (separated by periods) that describe that morpheme’s function and/or meaning. Labels in all caps (e.g. PAST) are grammatical glosses; lowercase labels (e.g. speak) are lexical glosses.
IGT examples may also use unsegmented transcriptions, as in the Uspanteko example in subsection 2.1 Pixabaj et al. (2007). In this case, words and their labels are aligned, but no explicit alignment between morpheme labels and the original morphemes themselves is present, and thus the segmentation of words into morphemes is unclear:
@@
IGT is highly valuable to documentation projects, and has also been used in the NLP context for the creation of morphological paradigms (Moeller et al., 2020), machine translation (Zhou et al., 2019), generating precision grammars (Bender et al., 2013), and other tools including POS taggers and dependency parsers (Georgi, 2016).
2.2 Challenges with Automatic Glossing
Effective glossing requires knowledge of the target language as well as linguistic understanding of morphological patterns. Furthermore, certain factors exist that make this task particularly difficult for automated systems.
Lack of segmented transcriptions: As in subsection 2.1, often only an unsegmented transcription is available. In these cases, an automated system cannot predict glosses morpheme-by-morpheme, but must instead learn to perform simultaneous segmentation and glossing.
Variance in glossing styles: While there are widely-used standards for glossing conventions, such as the Leipzig glossing rules Comrie et al. (2008), exact glossing styles may vary widely from documenter to documenter Chelliah et al. (2021). Documentation projects may have differences in label spelling (e.g. SING/SG/S to denote singular), formatting (e.g. underscores/periods for multi-word lexical glosses), and language specific labels (e.g. a linguist may want to specify different levels of an intensifier marker), among others.
Extremely low-data settings: When IGT is used for language documentation, the target languages of interest are likely those with very small amounts of digital data to begin with.
Given these challenges, we want our model to be easily adaptable to different languages with varying amounts of training data, as well as able to handle both segmented and unsegmented inputs.
3 Dataset
@@ 3.1 Data Sources
The Online Dictionary of Interlinear Text (ODIN, Lewis and Xia 2010) is a large dataset of 158k IGT examples representing 1496 languages, compiled by scraping IGT from linguistics documents on the internet. We use the preprocessed version of ODIN by He et al. (2023), which discards languages with less than five IGT samples, resulting in 84k unique glossed sentences across 936 languages.
@@ 3.2 Preprocessing
In total we have 250k unique IGT instances in 1.8k languages. We use Hugging Face Datasets to store our corpus in a widely-used, publicly accessible format. Each row retains a unique ID and reference to its source, so examples can be easily cross-referenced.
If datasets explicitly indicate whether an IGT line is segmented, we use this value. Otherwise, we determine segmentation by checking if a line has any morpheme boundary markers (the hyphen “-”). For segmented words, we remove the segmentation markers to create an additional unsegmented version of the same example, for a total of 451k examples (206k segmented).
@@ 3.4 Target Languages
We select data from the SIGMORPHON Shared Task to use for evaluation and testing, as this data is highly consistent and has been evaluated on prior models. We use the same splits as the shared task (Table 2) and designate three languagues–Arapaho, Tsez, and Uspanteko–as in-domain languages, which are included in the pretraining corpus. We use Gitksan, Lezgi, Natugu, and Nyangbo as out-of-domain languages, which do not occur in the pretraining corpus, but are used to finetune models. All languages except Nyangbo include translations.
@@
Arapaho (arp) 78,264 78,264 9,784 9,784
Tsez (ddo) 7116 7116 890 890
Uspanteko (usp) 19,584 19,584 464 1,266
@@
Gitksan (git) - 163 84 74
Lezgi (lez) - 1,411 176 174
Natugu (ntu) - 1,582 198 198
Nyangbo (nyb) - 4,200 526 526
Table 2: Number of total (segmented and unsegmented) pretrain (for in-domain languages), train, evaluation, and test examples for the target languages.
@@ 4 GlossLM Model
Using our IGT dataset, we aimed to create a single multilingual pretrained model that could be easily adapted to documentation projects, both for known languages and unseen ones.
@@ 4.1 Architecture
We base GlossLM off of ByT5, a multilingual pretrained model using the T5 architecture Raffel et al. (2020). ByT5 operates on byte-level inputs, as opposed to word or subword tokens, making it easily adaptable to different scripts and unseen languages. We use the ByT5-base model (582M parameters), pretrained on the mC4 dataset Xue et al. (2021). We did not experiment with pretraining a randomly initialized model, as pretraining runs are expensive and we predict that the pretrained non-IGT base model serves as a better initialization.
@@ 4.2 IGT Pretraining
We continually pretrain ByT5 on two versions of the train split of our dataset. The first (GlossLMall) uses all the training data in our dataset. Our second model (GlossLMunseg) excludes any segmented data in any of our test languages, simulating scenarios where a linguist does not have access to a corpus of segmented glossed data in their target language. Then, for each version of GlossLM, we finetune on the target language of interest, again using all data in the target language for GlossLMall and only unsegmented data in the target language for GlossLMunseg.
As ByT5 is based on the T5 model, we cast the glossing task into a text-to-text format. We train the model using examples formatted with the following prompt:
@@
Models are trained to output the gloss line following the above prompt input. Including translations has been shown to help in predicting glosses (Zhao et al., 2020; Ginn et al., 2023). For some data, a translation was not available ($\approx 3\%$ of the training data), in which case the translation line is omitted. We pretrain models using the hyperparameters given in Appendix C. We did not conduct hyperparameter search, only tuning the batch size, to fit in our GPUs, and epochs and early stopping, to ensure convergence.
5 Experiments
We evaluate the effectiveness of our pretrained model on automatic generation of IGT across the seven evaluation languages described in §3.4.
5.1 Experimental Settings
We compare models under two settings.
Segmented
In the segmented setting, words in the input transcriptions are divided into morphemes with segmentation boundaries. Models must learn to classify each morpheme with an appropriate gloss, making this an easier setting (with SOTA models achieving 80-90% accuracy).
Unsegmented
In the unsegmented setting, words in the input transcriptions are not segmented (but glosses are), requiring the model to jointly learn segmentation and glossing. Generally, this setting is far more difficult, with SOTA models achieving as low as 11% accuracy on small datasets.
5.2 GlossLM Experiments
We evaluate the GlossLM model in two variations. First, we directly evaluate the pretrained model on the in-domain languages without any additional finetuning, in order to analyze whether the model can retain high performance across individual languages despite the highly multilingual pretraining corpus. We refer to these models as GlossLMpre.
We then finetune the GlossLM pretrained model on the training dataset for each language individually, and label these runs as GlossLMft. We do so for both the in-domain languages, to focus the model on the target language, as well as the out-of-domain languages, allowing us to study the model’s adaptation to unseen languages. Finetuning used the same parameters as pretraining, but with 100 training epochs and early stopping (patience 3, start epoch 15) [^3] For Gitksan, due to the size of the training set, we set max epochs to 300 and patience to 15., and took anywhere from 20 minutes to around 4 hours for each language. Inference uses beam search with $n=3$ beams.
5.3 Baseline Experiments
We compare the GlossLM models with six baseline systems which include the state-of-the-art from prior work.
Top-choice
selects the most frequent label associated with each morpheme/word in the training data, and assigns “???” to unseen morphemes.
Token-class
treats the glossing task as a token classification problem, where the output vocabulary consists of the IGT morpheme or word-level labels. Each target language’s data is used to train a language-specifc Token-class model, which uses the RoBERTa architecture with default hyperparameters without any additional pretraining Liu et al. (2019). This was used as the baseline model for the SIGMORPHON 2023 Shared Task on interlinear glossing Ginn (2023).
Tü-CL Girrbach (2023b)
uses straight-through gradient estimation (Bengio et al., 2013) to induce latent, discrete segmentations of input texts, and predicted glosses using a multilayer perceptron.
CRF Okabe and Yvon (2023b)
is a hybrid CRF-neural system that is composed of an unsupervised neural alignment system (SimAlign, Jalili Sabet et al. 2020) to make use of lexemes in translations and a CRF sequence labeling system (Lost, Lavergne et al. 2011).
SigMoreFun (SMF) He et al. (2023)
also continually pretrain ByT5, though on a smaller amount of data (using the ODIN corpus). Data from all target languages were combined and shuffled for finetuning.
ByT5all and ByT5unseg
are ByT5 models Xue et al. (2022) finetuned on the target language data without any multilingual continual pretraining on IGT. We otherwise replicate the finetuning procedure described in subsection 5.2.
5.4 Evaluation
For evaluating predictions, we strip punctuation (except for within glosses). [^4] Because of this post-processing, our results for baselines are slightly different than what original sources report. We evaluate morpheme accuracy, which counts the number of correct morpheme glosses in the correct position. Hence, if a gloss is inserted or deleted, the following glosses will be incorrect. We also evaluate word accuracy, which counts the number of entire correct word glosses.
Model Morpheme accuracy / Word accuracy (Segmented)
In-domain Out-of-domain
arp ddo usp Avg git lez ntu nyb Avg
Model Morpheme accuracy / Word accuracy (Unsegmented)
In-domain Out-of-domain
arp ddo usp Avg git lez ntu nyb Avg
GlossLMunseg, ft 82.1 / 81.5 83.6 / 87.3 78.6 / 81.0 81.4 / 83.3 10.1 / 28.4 57.3 / 64.9 62.8 / 78.9 87.4 / 86.2 56.3 / 65.0

### v2

# GlossLM: A Massively Multilingual Corpus and Pretrained Model for Interlinear Glossed Text

@@ Abstract
Language documentation projects often involve the creation of annotated text in a format such as interlinear glossed text (IGT), which captures fine-grained morphosyntactic analyses in a morpheme-by-morpheme format. However, there are few existing resources providing large amounts of standardized, easily accessible IGT data, limiting their applicability to linguistic research, and making it difficult to use such data in NLP modeling.
We compile the largest existing corpus of IGT data from a variety of sources, covering over 450k examples across 1.8k languages, to enable research on crosslingual transfer and IGT generation. We normalize much of our data to follow a standard set of labels across languages.
Furthermore, we explore the task of automatically generating IGT in order to aid documentation projects. As many languages lack sufficient monolingual data, we pretrain a large multilingual model on our corpus. We demonstrate the utility of this model by finetuning it on monolingual corpora, outperforming SOTA models by up to 6.6%. We will make our pretrained model and dataset available through Hugging Face, as well as provide access through a web interface for use in language documentation efforts.
@@
GlossLM: A Massively Multilingual Corpus and Pretrained Model for Interlinear Glossed Text
Michael Ginn*1  Lindia Tjuatja*2  Taiqi He2  Enora Rice1 Graham Neubig2Alexis Palmer1Lori Levin2 1University of Colorado Boulder  2Carnegie Mellon University michael.ginn@colorado.edulindiat@andrew.cmu.edu * Equal contribution
@@ 1 Introduction
With nearly half of the world’s 7,000 languages considered endangered, communities of minoritized language speakers are working to preserve and revitalize their languages (Seifart et al., 2018). These efforts often involve collection, analysis, and annotation of linguistic data. Annotated text can be used in the creation of reference materials (such as dictionaries and grammars) as well as to develop language technologies including searchable digital text Blokland et al. (2019); Rijhwani et al. (2023) and computer-assisted educational tools Uibo et al. (2017); Chaudhary et al. (2023).
@@
Interlinear glossed text (IGT) is a widespread format in language documentation for linguistic annotation. IGT is a multi-line data format (see Figure 1) which includes (1) a transcription of speech in the language, (2) an aligned morpheme-by-morpheme description, and oftentimes (3) a free translation. IGT can be used to illustrate morphosyntactic features of languages that other researchers may not be familiar with, and it is a popular format for examples in linguistics papers and textbooks. It also serves as a resource in the NLP context for the creation of morphological paradigms (Moeller et al., 2020), machine translation (Zhou et al., 2019), generating precision grammars (Bender et al., 2013), and other tools including POS taggers and dependency parsers (Georgi, 2016).
Compiling IGT Data
Though IGT often follows a common glossing format, gloss conventions vary wildly. Furthermore, IGT data is rarely compiled into large, standardized corpora, often existing as scattered examples in research papers. To address this, we compile the largest corpus of digitized IGT from various existing sources, with over 450k examples in 1.8k languages (§3). We explore methods to normalize this data (§4)., standardizing over 80% of the grammatical glosses in the corpus to follow the UniMorph (Sylak-Glassman, 2016). We will release our corpora for future NLP, linguistics, and language documentation research.
Automating IGT Generation
Futhermore, creating new IGT corpora is difficult and time-consuming, requiring documenters to perform linguistic analysis and extensive documentation simultaneously. Research has found that computational tools can help accelerate annotation and overcome this bottleneck Palmer et al. (2009) by predicting the gloss line of IGT given a transcription.
The majority of prior work on automatic glossing focuses on training monolingual models that can predict IGT for a single language (Moeller and Hulden, 2018; McMillan-Major, 2020; Zhao et al., 2020), however, these models can struggle when data is limited and require dedicated effort to train and deploy.
We aim to overcome the monolingual data bottleneck by creating a multilingual pretrained glossing model that can be adapted to specific languages and gloss formats. We continually pretrain a model on our corpus, finding that pretrained multilingual model retains high performance across languages. We then finetune the pretrained model on monolingual data, including languages that do not appear in the pretraining corpus. We achieve a new SOTA performance on five out of seven languages, and demonstrate clear improvements for low-resource language settings over an equivalent finetuned model without our continual pretraining (§7).
@@ 2 Interlinear Glossed Text (IGT)
2.1 Format
Interlinear glossed text is a structured data format which presents text in a language being studied along with morphological glosses—aligned labels that indicate each morpheme’s meaning and/or grammatical function. Often, a free translation in a widely-spoken language is included as well. An IGT example for Arapaho is given in subsection 2.1 Cowell (2020), with glosses and translations in English.
@@ 2.1 Format
This example is segmented, with morphemes separated by dashes. Each morpheme in the Arapaho sentence (e.g. tih) is directly aligned with a gloss (e.g. when.Past) that describes the morpheme’s function and/or meaning. Labels in all caps (e.g. Past) are grammatical glosses; lowercase labels (e.g. speak) are lexical glosses. Periods are used for fusional morphemes, which carry several morphological or lexical functions in a single morpheme.
IGT examples may instead use unsegmented transcriptions, as in the Uspanteko example in subsection 2.1 Pixabaj et al. (2007).
@@
Here, words and their labels are aligned, but no explicit alignment between morpheme glosses and individual morphemes is provided, and thus the segmentation of words into morphemes is unclear.
2.2 Challenges with Interlinear Glossing
Effective glossing requires expert knowledge of the target language and linguistic understanding of morphological patterns. Furthermore, certain factors exist that make this task particularly difficult for automated systems. Often, transcriptions are not segmented into morphemes, and systems must perform simultaneous segmentation and glossing.
Glossing conventions and formats may vary widely from documenter to documenter Chelliah et al. (2021), with differences in label spelling (e.g. Sing/Sg/S to denote singular), formatting and punctuation, and language-specific labels. Furthermore, nearly all languages have very little digitized IGT data, posing difficulty to automated systems and human annotators alike. Finally, even when automated systems have been created, practical deployment remains an additional challenge for documenters.
3 GlossLM Corpus
@@ 3.1 Data Sources
The Online Dictionary of Interlinear Text (ODIN, Lewis and Xia 2010) is a large dataset of 158k IGT examples representing 1496 languages, compiled by scraping IGT from linguistics documents on the internet. We use the preprocessed version of ODIN by He et al. (2023), which discards languages with fewer than five IGT samples, resulting in 84k unique glossed sentences across 936 languages.
@@ 3.2 Preprocessing
In total we have 250k unique IGT instances in 1.8k languages. If datasets explicitly indicate whether an IGT line is segmented, we use this value. Otherwise, we determine segmentation by checking if a line has any morpheme boundary markers (the hyphen “-”). For segmented words, we remove the segmentation markers to create an additional unsegmented version of the same example, for a total of 451k examples (206k segmented).
@@ 3.3 Language Coverage
4 Normalizing Gloss Labels
4.1 Motivation
As our data comes from a variety of sources, spanning many languages and documentation projects, there is a great amount of diversity of the morphological glosses used. This includes cases where several different labels are used to indicate the same feature (e.g. Sing, Sg, or S for singular), as well as formatting differences such as the usage of periods (e.g. 1Sg vs 1.Sg).
We explore the value and feasibility of normalizing glosses to a single standardized format. On one hand, normalizing glosses may make it easier to train models that utilize crosslingual information through shared gloss labels, but on the other hand, it can be difficult to select a single schema that preserves the original intent of all annotators.
We split gloss lines by period and count the number and frequency of unique grammatical gloss labels across our corpus (focusing on the all-caps functional glosses, not stems) and visualize the distribution in Figure 2.
Figure 2: Distribution of unique glosses across all languages.
There are 11,493 unique glosses which roughly form a Zipfian distribution (Zipf, 1945). The most common glosses, unsurprisingly, are labels such as Pl (plural, 52,488 instances), 3Sg (3rd-person singular, 39,147), and Past (36,124), which occur broadly across many languages.
Normalizing all unique glosses would be a monumental task with uncertain benefits. However, we observe that the 200 most common gloss types account for 82.7% of glosses in our dataset. We focus on normalizing these glosses: e.g. all instances of Past and Pst, which are both in the top 200, should be normalized to the same label. We note that there are other aspects of the data that vary (e.g. periods vs. underscores for multi-word glosses, representing non-concatenative morphology), as addressed in Mortensen et al. (2023). Future work could potentially focus on the benefits of these aspects of normalization on training.
4.2 Methodology
We select the UniMorph schema of Sylak-Glassman (2016) as our standardized set. While no single set of labels captures the intricacies of all of the world’s languages, UniMorph is widely used and has coverage for many common features.
The two lead authors of this paper jointly created a mapping from the labels in our dataset to UniMorph labels. While many mappings were obvious (or already compatible with UniMorph), others posed a myriad of issues. For glosses primarily used for a single language, we consulted the original source dataset to determine the meaning.
Ambiguous labels
Several labels were ambiguous, corresponding to one of several UniMorph glosses depending on the language and annotator. For example, the label S (appearing 20,855 times) is used for singular, subject, and noun/sustantivo (at least) in our dataset. In order to map these, we would need to analyze on a sentence-by-sentence basis, which was not feasible; thus, we left these labels as-is.
Glosses not in UniMorph
UniMorph primarily focuses on common crosslingual inflectional features, and does not cover the full extent of the morphological systems of the world. We observed 64 of 200 (32%) gloss labels with no clear UniMorph equivalent, including demonstratives (Dem, 15,585 instances), obliques (Obl, 13,639), and clitics (Cl, 7,453). In many cases, there is a related UniMorph gloss that is more general or more specific; for example preterite (Pret, 1,986) could be mapped to simple past (Pst). However, this would be an imprecise mapping, and could be confusing to a linguist of the particular language, so we again elect to leave these glosses unmapped.
We use our mapping to normalize the dataset and will make the normalized version available in addition to the original.
4.3 Use in Future Research
We believe our dataset can potentially be useful across NLP research, linguistic research, and language documentation. NLP researchers benefit from a single, easily-accessible dataset covering many languages, which can be used for future research on interlinear glossing, morpheme analysis, segmentation, and translation.
Linguistics researchers will be able to use the dataset to easily search for phenomena across languages, particular with the normalized version of the dataset. For example, a linguist could find examples of sentences demonstrating the ergative/absolutive distinction. They could further refine this analysis by narrowing the results to related languages, using the glottocodes in our dataset.
As another example, if a linguist wished to determine how prior researchers have annotated examples in a particular language, they would previously have to search across research papers, textbooks, and small corpora. With our dataset, it is trivial to pull up all of the examples in a particular language, potentially compiled from many sources.
5 Automatic IGT Generation
Next, we evaluate the applicability of our dataset to the NLP task of automatic gloss prediction. We select the IGT data from the SIGMORPHON Shared Task on Interlinear Glossing (Ginn et al., 2023) to use for evaluation and testing, as this data consistently adheres to a set of glossing conventions and has been evaluated on prior models.
@@ 5.1 Target Languages
We reuse the train/eval/test splits for the seven languages from the SIGMORPHON Shared Task (Table 2). We designate three languagues—Arapaho, Tsez, and Uspanteko—as in-domain languages, which are included in the GlossLM corpus. We use Gitksan, Lezgi, Natugu, and Nyangbo as out-of-domain languages, which are omitted from the corpus. All languages except Nyangbo include translations.
The shared task included two distinct settings:
  • In the open track, the transcription lines were segmented into morphemes. This becomes a token classification task, and tends to be far easier, with SOTA models achieving 80-90% accuracy (and even a naïve method that simply selects the most common gloss for each morpheme was very effective).
  • In the closed track, transcription lines were not segmented. This setting is far more challenging, as models must jointly learn to segment words and predict glosses, and the best models achieved as low as 11% accuracy on small datasets. On the other hand, these models have the potential to be more valuable to documentation projects, where segmented text may not be available.
In our experiments, we focus on the unsegmented setting (closed track). However, the segmented data is also included in the GlossLM corpus, and could easily be used in future research.
@@
Arapaho (arp) 39,132 39,132 4,892 4,892
Tsez (ddo) 3,558 3,558 445 445
Uspanteko (usp) 9,774 9,774 232 633
@@
Gitksan (git) - 74 42 37
Lezgi (lez) - 705 88 87
Natugu (ntu) - 791 99 99
Nyangbo (nyb) - 2,100 263 263
Table 2: Number of total (unsegmented) pretrain (for in-domain languages), train, evaluation, and test examples for the target languages.
5.2 Evaluation Metrics
For evaluating predictions, we strip punctuation (except for within glosses). [^2] Because of this post-processing, our results for baselines are slightly different than what original sources report. We evaluate morpheme accuracy, which counts the number of correct morpheme glosses in the correct position. Hence, if a gloss is incorrectly inserted or deleted, the subsequent glosses will be incorrect. We also evaluate word accuracy, which counts the number of entire correct word glosses.
@@ 6 GlossLM Model
Using our IGT corpus described in section 3, we train a single multilingual pretrained model for the glossing task that can be easily adapted to documentation projects, for both seen languages and unseen ones.
@@ 6.1 Architecture
We use the ByT5 model, a multilingual pretrained model using the T5 architecture Raffel et al. (2020). ByT5 operates on byte-level inputs, as opposed to word or subword tokens, making it easily adaptable to different scripts and unseen languages. We use the ByT5-base model (582M parameters), pretrained on the mC4 dataset Xue et al. (2021). We did not experiment with pretraining a randomly initialized model, as pretraining runs are expensive and we predict that the pretrained non-IGT base model serves as a better initialization.
@@ 6.2 IGT Pretraining
We continually pretrain the ByT5 model on the GlossLM corpus described in section 3. As we are evaluating on unsegmented IGT data, we omit segmented data for the evaluation languages from the pretraining corpus. [^3] However, our corpus includes segmented IGT examples in other languages, which we do not evaluate. We structure the glossing task as a text-to-text problem, training the model with examples formatted with the following prompt:
@@
Models are trained to output the gloss line following the above prompt input. We include translations, which has been shown to provide benefits in gloss prediction (Zhao et al., 2020; Ginn et al., 2023). For some data, a translation was not available ($\approx 3\%$ of the training data), in which case the translation line is omitted. We pretrain models using the hyperparameters given in Appendix C. We did not conduct hyperparameter search, only tuning the batch size, to fit in our GPUs, and epochs and early stopping, to ensure convergence.
6.3 Performance of Pretrained Model
When training massively multilingual models, performance on individual languages can sometimes degrade in what is dubbed the "curse of multilinguality" (Conneau et al., 2020; Chang et al., 2023). To investigate this issue, we evaluate our pretrained model on the in-domain languages without any additional finetuning.
We compare the performance of our pretrained model to the current SOTA, which is the second system from Girrbach (2023a), as shown in Figure 3. We find that the model outperforms the SOTA across all three in-domain languages. This result gives little evidence to believe our model suffers from the curse of multilinguality, as it retains good performance across several languages. [^4] These languages do make up large fractions of our pretraining corpus, so the model will almost certainly underperform on underrepresented languages.
Figure 3: Comparison of our pretrained model and the SOTA Girrbach (2023a) for in-domain languages on unsegmented data. Our model outperforms on all three languages.
Our pretrained model can be used for automated glossing across several languages, without needing to train and serve separate monolingual models. This could be valuable to real-world documentation projects, as we can serve a single pretrained model that can be used across projects, significantly reducing the barrier to using an automated system.